Two people are attempting to push a 1000 kg car on a horizontal plan for 100 meters. Assume they are pushing on the car at a 20° angle upward with a force of 500N, how much work is being done?


1
Expert's answer
2021-06-03T09:19:16-0400
"W=Fs\cos{20}\\W=(500)(100)\cos{20}=47\ kJ"